# Generalized Linear Bandits with Limited Adaptivity

### v3

@@ 3 B-GLinCB
In this section, we present B-GLinCB (Algorithm 1) that solves Problem 1 described in Section 2, which enforces constraints of limited adaptivity setting M1. Given limited adaptivity budget $M\in\mathbb{N}$, our algorithm first computes the batch length for each of the $M$ batches (i.e., rounds where the policy remains constant). We build upon the batch length construction in Gao et al. (2019); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.
Batch Lengths: For each batch $k\in[M]$, let ${\mathcal{T}}_{k}$ denote all the consecutive rounds within the $k^{th}$ batch. We will refer to the first batch ${\mathcal{T}}_{1}$ as the warm-up batch. The batch lengths $\tau_{k}\coloneqq|{\mathcal{T}}_{k}|$, $k\in[M]$ are calculated as follows:
During the warm-up batch (Lines 2, 4), the algorithm follows the G-optimal design policy, $\pi_{G}$. At the end of the warm-up batch (Line 5), the algorithm computes the Maximum Likelihood Estimate (MLE), $\widehat{\theta}_{w}$, of $\theta^{*}$ [^4] In case the MLE lies outside the set $\{\theta^{*}:\left\lVert\theta^{*}\right\rVert\leq S\}$, we follow the projection step detailed in Appendix E., and design matrix $\mathbf{V}\coloneqq\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}$, with parameter $\lambda=20Rd\log{T}$.
Now, for each batch $k\geq 2$ and every round $t\in{\mathcal{T}}_{k}$, the algorithm updates ${\mathcal{X}}_{t}$ by eliminating arms from it using the confidence bounds (see Equation (5)) computed in the previous batches (Line 12). The algorithm next computes $\widetilde{{\mathcal{X}}}_{t}$, a scaled version of ${\mathcal{X}}_{t}$, as follows
$\widetilde{{\mathcal{X}}}_{t}\coloneqq\left\{\sqrt{{\dot{\mu}(\langle x,\widehat{\theta}_{w}\rangle)}/{\beta(x)}}\ \ \ x:\text{ for all }\ x\in{\mathcal{X}}_{t}\right\}.$ (2)
Finally, we use the distributional optimal design policy $\pi_{k}$, on the scaled arm set $\widetilde{{\mathcal{X}}}_{t}$, to sample the next arm (Line 13). At the end of every batch, we equally divide the batch ${\mathcal{T}}_{k}$ into two sets ${\mathcal{A}}$ and ${\mathcal{B}}$. We use samples from ${\mathcal{A}}$ to compute the estimator $\widehat{\theta}_{k}$ and the scaled design matrix ${\mathbf{H}}_{k}$. The rounds in ${\mathcal{B}}$ are used to compute $\pi_{k+1}$, the distributional optimal design policy for the next batch. It is important to note, while the policy $\pi_{k}$ is utilized in each round (Line 13) to draw arms, it is updated (to $\pi_{k+1}$) only at the end of the batch. Hence, conforming to setting M1, the algorithm updates the selection policy at $M$ rounds that were decided upfront.
Remark 3.1.
@@
Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$, where
@@
The $\kappa$-dependent batch construction is a crucial feature of our algorithm, enabling effective estimation of $\dot{\mu}{\langle\theta,x\rangle}$ at the end of the first batch. Since the first batch incurs regret linear in its length, achieving a $\kappa$-independent guarantee requires the first batch to be $o(\sqrt{T})$. We demonstrate that choosing $\tau_{1}=O(T^{\frac{1}{3}})$ is sufficient for this purpose (see Appendix A).
@@ 4 RS-GLinCB
15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{w}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
@@
When Switching Criterion I is not triggered, the algorithm first checks (Line 9) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered). If Switching Criterion II is triggered at round $t$, then in Line 10, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$. Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^5] This optimization problem is non-convex. However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$,
$\min_{\theta}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}\right\rVert_{{\mathbf{H}}(\theta)}\text{s.t. }\left\lVert\theta-\widehat{\theta}_{w}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (6)
Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$. After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$. Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$. For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}\sqrt{d\log{({T}/{\delta})}}$, and plays the arm $x_{t}$ with the highest index (Line 14). After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 15) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$. With this, the round $t$ ends and the algorithm moves to the next round. Next, in Lemma 4.1 and Theorem 4.2 we present the guarantees on number of policy updates and regret, respectively, for RS-GLinCB. Detailed proofs for both are provided in Appendix B.
@@
Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting. Unlike existing literature (Faury et al., 2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.4 and B.5). We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.3), hence incurring a small regret in these rounds.
@@
Unlike (Abbasi-Yadkori et al., 2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$). The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret. This modification is a crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
@@
The feasible set for the optimization stated in 6 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability. Deviating from existing literature on GLM Bandits which projects the estimate into the ball set of radius $S$ ($\{\theta:\left\lVert\theta\right\rVert\leq S\}$), our projection step leads to tighter regret guarantees; notably, the leading $\sqrt{T}$ term is free of parameters $S$ (and $R$). This resolves the conjecture made in Lee et al. (2023) regarding the possibility of obtaining $S$-free regret in the $\sqrt{T}$ term in logistic bandits.
@@ 5 Experiments
Figure 1: Top: Cumulative Regret vs. number of rounds for Logistic (left) and Probit (right) reward models. Bottom: (left) Execution times of ECOLog and RS-GLinCB for different values of $\kappa$ (low $\kappa=9.3$ and high $\kappa=141.6$) for Logistic rewards. (right) Execution times of GLOC and RS-GLinCB for different values of $\kappa$ (low $\kappa=17.6$ and high $\kappa=202.3$) Probit rewards.
We tested the practicality of our algorithm RS-GLinCB against various baselines for logistic and generalized linear bandits. For these experiments, we adjusted the warm-up threshold constant in RS-GLinCB to $0.01$ and used data from both warm-up and non-warm-up rounds to estimate $\widetilde{\theta}$. These modifications do not affect the overall efficiency as $\widetilde{\theta}$ is calculated only $O(\log(T))$ times. The experiment code is available at https://github.com/nirjhar-das/GLBandit_Limited_Adaptivity.
Logistic. We compared RS-GLinCB against ECOLog (Faury et al., 2022) and GLOC (Jun et al., 2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting. The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$. Arms were sampled uniformly from the $d$-dimensional unit ball. We ran simulations for $T=20,000$ rounds, repeating them 10 times. RS-GLinCB showed the smallest regret with a flattened regret curve, as seen in Fig. 1 (top-left).
Probit. For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB (Filippi et al., 2010). The dimension was set to $d=5$ and number of arms per round to $K=20$. $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=3$. Arm features were generated similarly as in the logistic bandit simulation. We ran simulations for $T=5,000$ rounds, repeating them 10 times. RS-GLinCB outperformed both baselines, as shown in Fig. 1 (top-right).
@@ 6 Conclusion and Future Work
The Contextual Bandit problem with GLM rewards is a ubiquitous framework for studying online decision-making with non-linear rewards. We study this problem with a focus on limited adaptivity. In particular, we design algorithms B-GLinCB and RS-GLinCB that obtain optimal regret guarantees for two prevalant limited adaptivity settings M1 and M2 respectively. A key feature of our guarantees are that their leading terms are independent of a instance dependent parameter $\kappa$ that captures non-linearity. To the best of our knowledge, our paper provides the first algorithm for CB algorithms for GLM rewards under limited adaptivity (and otherwise) that achieves $\kappa$-independent regret. The regret guarantee of RS-GLinCB, not only aligns with the best-known guarantees for Logistic Bandits but enhances them by removing the dependence on $S$ (upper bound on $\left\lVert\theta^{*}\right\rVert$) in the leading term of the regret and therefore resolves a conjecture in Lee et al. (2023). The batch learning algorithm B-GLinCB, for $M=\Omega(\log{(\log{T})})$, achieves a regret of $\widetilde{O}\left(dRS\left(\sqrt{{d/{\widehat{\kappa}}}}\land\sqrt{1/\kappa^{*}}\right)\sqrt{T}\right)$. We believe that the dependence on $d$ along with the ${\widehat{\kappa}}$ term is not tight and improving the dependence is a relevant direction for future work.
Appendix for “”

### v4

@@ 3 B-GLinCB
In this section, we present B-GLinCB (Algorithm 1) that solves Problem 1 described in Section 2, which enforces constraints of limited adaptivity setting M1. Given limited adaptivity budget $M\in\mathbb{N}$, our algorithm first computes the batch length for each of the $M$ batches (i.e., determine rounds where the policy remains constant). We build upon the batch length construction in Gao et al. (2019); however, the first batch is chosen to be $\kappa$ dependent which crucially helps in removing $\kappa$ from the leading term in the regret.  [^5] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm.
Batch Lengths: For each batch $k\in[M]$, let ${\mathcal{T}}_{k}$ denote all the rounds within the $k^{th}$ batch. We will refer to the first batch ${\mathcal{T}}_{1}$ as the warm-up batch. The batch lengths $\tau_{k}\coloneqq|{\mathcal{T}}_{k}|$, $k\in[M]$ are calculated as follows:
During the warm-up batch (Lines 2, 4), the algorithm follows the G-optimal design policy, $\pi_{G}$. At the end of the warm-up batch (Line 5), the algorithm computes the Maximum Likelihood Estimate (MLE), $\widehat{\theta}_{w}$, of $\theta^{*}$ [^7] In case the MLE lies outside the set $\{\theta^{*}:\left\lVert\theta^{*}\right\rVert\leq S\}$, we apply the projection step detailed in Appendix E., and design matrix $\mathbf{V}\coloneqq\sum_{t\in{\mathcal{T}}_{1}}x_{t}x_{t}^{\mathsf{T}}+\lambda{\mathbf{I}}$, with parameter $\lambda=20Rd\log{T}$.
Now, for each batch $k\geq 2$ and every round $t\in{\mathcal{T}}_{k}$, the algorithm updates ${\mathcal{X}}_{t}$ by eliminating arms from it using the confidence bounds (see Equation (7)) computed in the previous batches (Line 12). The algorithm next computes $\widetilde{{\mathcal{X}}}_{t}$, a scaled version of ${\mathcal{X}}_{t}$, as follows, with $\beta(x)$ define in equation  (5),
$\widetilde{{\mathcal{X}}}_{t}\coloneqq\left\{\sqrt{{\dot{\mu}(\langle x,\widehat{\theta}_{w}\rangle)}/{\beta(x)}}\ \ \ x:\ x\in{\mathcal{X}}_{t}\right\}.$ (4)
Finally, we use the distributional optimal design policy $\pi_{k}$, on the scaled arm set $\widetilde{{\mathcal{X}}}_{t}$, to sample the next arm (Line 13). At the end of every batch, we equally divide the batch ${\mathcal{T}}_{k}$ into two sets ${\mathcal{A}}$ and ${\mathcal{B}}$. We use samples from ${\mathcal{A}}$ to compute the estimator $\widehat{\theta}_{k}$ and the scaled design matrix ${\mathbf{H}}_{k}$. The rounds in ${\mathcal{B}}$ are used to compute $\pi_{k+1}$, the distributional optimal design policy for the next batch. It is important to note while the policy $\pi_{k}$ is utilized in each round (Line 13) to draw arms, it is updated (to $\pi_{k+1}$) only at the end of the batch. Hence, conforming to setting M1, the algorithm updates the selection policy at $M$ rounds that were decided upfront.
Remark 3.1.
@@
Algorithm 1 (B-GLinCB) incurs regret $\textsc{R}_{T}\leq(\textsc{R}_{1}+\textsc{R}_{2})\log{\log{T}}$  [^8] Note that $\textsc{R}_{T}$ is expected regret. See the 2., where
@@
The $\kappa$-dependent batch construction is a crucial feature of our algorithm, enabling effective estimation of $\dot{\mu}{\left(\langle x,\theta^{*}\rangle\right)}$ at the end of the first batch. Since the first batch incurs regret linear in its length, achieving a $\kappa$-independent guarantee requires the first batch to be $o(\sqrt{T})$. We demonstrate that choosing $\tau_{1}=O(T^{\frac{1}{3}})$ is sufficient for this purpose (see Appendix A).
@@ 4 RS-GLinCB
15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{o}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
@@
When Switching Criterion I is not triggered, the algorithm first checks (Line 9) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered). If Switching Criterion II is triggered at round $t$, then in Line 10, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$. Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^9] This optimization problem is non-convex. However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
$\min_{\theta}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}\right\rVert_{{\mathbf{H}}(\theta)}\text{s.t. }\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (8)
Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$. After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ [^10] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm.. Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$. For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}\sqrt{d\log{({T}/{\delta})}}$, and plays the arm $x_{t}$ with the highest index (Line 14). After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 15) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$. With this, the round $t$ ends and the algorithm moves to the next round. Next, in Lemma 4.1 and Theorem 4.2 we present the guarantees on number of policy updates and regret, respectively, for RS-GLinCB. Detailed proofs for both are provided in Appendix B.
@@
Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting. Unlike existing literature Faury et al. (2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.3 and B.4). We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.11), hence incurring a small regret in these rounds.
@@
Unlike Abbasi-Yadkori et al. (2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$). The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret. This modification is crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
@@
The feasible set for the optimization stated in 8 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability. Deviating from existing literature on GLM Bandits which projects the estimate into the ball set of radius $S$ ($\{\theta:\left\lVert\theta\right\rVert\leq S\}$), our projection step leads to tighter regret guarantees; notably, the leading $\sqrt{T}$ term is free of parameters $S$ (and $R$). This resolves the conjecture made in Lee et al. (2024) regarding the possibility of obtaining $S$-free regret in the $\sqrt{T}$ term in logistic bandits.
Remark 4.6.
The regret guarantees of the logistic bandit algorithms in Abeille et al. (2021); Lee et al. (2024) have a second-order term that is minimum of an arm-geometry dependent quantity (see Theorem 3 of Lee et al. (2024)) and a $\kappa$-dependent term similar to our regret guarantee. Although our analysis is not able to accommodate this arm-geometry dependent quantity, we underscore that our algorithm is computationally efficient while the above works are not. In fact, to the best of our knowledge, the other known efficient algorithms for logistic bandits Faury et al. (2022); Zhang and Sugiyama (2023) also do not achieve the arm-geometry dependent regret term. It can be interesting to design an efficient algorithm that is able to achieve the same guarantees in the second-order regret term as in Abeille et al. (2021); Lee et al. (2024).
@@ 5 Experiments
Figure 1: Top: Cumulative Regret vs. number of rounds for Logistic (left) and Probit (right) reward models. Bottom: (left) Execution times of ECOLog and RS-GLinCB for different values of $\kappa$ (low $\kappa=9.3$ and high $\kappa=141.6$) for Logistic rewards. (right) Execution times of GLOC and RS-GLinCB for different values of $\kappa$ (low $\kappa=17.6$ and high $\kappa=202.3$) for Probit rewards.
We tested the practicality of our algorithm RS-GLinCB against various baselines for logistic and generalized linear bandits. For these experiments, we adjusted the Switching Criterion I threshold constant in RS-GLinCB to $0.01$ and used data from both Switching Criteria (I and II) rounds to estimate $\widetilde{\theta}$. These modifications do not affect the overall efficiency as $\widetilde{\theta}$ is calculated only $O(\log(T))$ times. The experiment code is available at https://github.com/nirjhar-das/GLBandit_Limited_Adaptivity.
Logistic. We compared RS-GLinCB against ECOLog Faury et al. (2022) and GLOC Jun et al. (2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting. The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$. Arms were sampled uniformly from the $d$-dimensional unit ball. We ran simulations for $T=20,000$ rounds, repeating them 10 times. RS-GLinCB showed the smallest regret with a flattened regret curve, as seen in Fig. 1 (top-left).
Probit. For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB Filippi et al. (2010). The dimension was set to $d=5$ and number of arms per round to $K=20$. $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=3$. Arm features were generated similarly as in the logistic bandit simulation. We ran simulations for $T=5,000$ rounds, repeating them 10 times. RS-GLinCB outperformed both baselines, as shown in Fig. 1 (top-right).
@@ 6 Conclusion and Future Work
The Contextual Bandit problem with GLM rewards is a ubiquitous framework for studying online decision-making with non-linear rewards. We study this problem with a focus on limited adaptivity. In particular, we design algorithms B-GLinCB and RS-GLinCB that obtain optimal regret guarantees for two prevalant limited adaptivity settings M1 and M2 respectively. A key feature of our guarantees are that their leading terms are independent of an instance dependent parameter $\kappa$ that captures non-linearity. To the best of our knowledge, our paper provides the first algorithms for the CB problem with GLM rewards under limited adaptivity (and otherwise) that achieve $\kappa$-independent regret. The regret guarantee of RS-GLinCB, not only aligns with the best-known guarantees for Logistic Bandits but enhances them by removing the dependence on $S$ (upper bound on $\left\lVert\theta^{*}\right\rVert$) in the leading term of the regret and therefore resolves a conjecture in Lee et al. (2024). The batch learning algorithm B-GLinCB, for $M=\Omega(\log{(\log{T})})$, achieves a regret of $\widetilde{O}\left(dRS\left(\sqrt{{d/{\widehat{\kappa}}}}\land\sqrt{1/\kappa^{*}}\right)\sqrt{T}\right)$. We believe that the dependence on $d$ along with the ${\widehat{\kappa}}$ term is not tight and improving the dependence is a relevant direction for future work.
Acknowledgments and Disclosure of Funding
Siddharth Barman gratefully acknowledges the support of the Walmart Center for Tech Excellence (CSR WMGT-23-0001) and a SERB Core research grant (CRG/2021/006165).
Appendix